# Near-Optimal Emission-Aware Online Ride Assignment Algorithm for Peak Demand Hours

### v1

@@
Ridesharing has experienced significant global growth over the past decade and is becoming integral to future transportation networks. These services offer alternative mobility options in many urban areas, promoting car-light or car-free lifestyles, with their market share rapidly expanding due to the convenience they offer. However, alongside these benefits, concerns have arisen about the environmental impact of ridesharing, particularly its contribution to carbon emissions. A major source of these emissions is “deadhead miles” that are driven without passengers between trips. This issue is especially pronounced during high-demand periods when the number of ride requests exceeds platform capacity, leading to longer deadhead miles and higher emissions. While reducing these unproductive miles can lower emissions, it may also result in longer wait times for passengers as they wait for a nearby driver, potentially diminishing the overall user experience. As ridesharing platforms continue to grow, solutions that reduce carbon emissions must strike a balance between environmental objectives and maintaining service quality, particularly the wait times of the rider.
In this paper, we propose LARA, an online algorithm for rider-to-driver assignment that dynamically adjusts the maximum allowed deadhead miles for drivers and assigns ride requests accordingly. While LARA can be applied under any conditions, it is particularly more effective during high-demand hours, aiming to reduce both carbon emissions and rider wait times. We prove that LARA achieves near-optimal performance in online settings compared to the optimal offline algorithm. Furthermore, we evaluate LARA using both synthetic and real-world datasets, demonstrating up to $34.2\%$ reduction in emissions and up to $42.9\%$ reduction in rider wait times compared to state-of-the-art algorithms. While recent studies have introduced the problem of emission-aware ride assignment, LARA is the first algorithm to provide both theoretical and empirical guarantees on performance.
@@ 1 Introduction
In 2022, the transportation sector accounted for 28.4% of total greenhouse gas emissions. This has been a consistent trend, as transportation has remained the leading source of emissions in the U.S. since 2017. The dominance of transportation in emissions is largely due to the widespread use of vehicles powered by gasoline and diesel, which are major sources of carbon dioxide and other greenhouse gases. As the population grows, economies develop, and urbanization continues, the demand for urban transport is expected to increase. This increase is likely to counteract any CO2 reductions from low- and zero-carbon technologies. By 2050, projections indicate that motorized urban mobility will nearly double, leading to a 26% rise in CO2 emissions from urban transport [14]. Given these environmental challenges, there is a growing need to investigate and develop transportation systems that focus on sustainable practices.
With growing populations and advancing technology, ridesharing services have transformed travel by offering on-demand mobility through individual or shared vehicles, leading to global adoption [35]. According to Statista, the global ridesharing market is expected to experience substantial growth, with revenue anticipated to reach \$$\$167.60$ billion by 2024 and \$$\$212.80$ billion by 2029 [31]. The number of users is also projected to increase, reaching 2.31 billion by 2029, with user penetration rising from 23.1% in 2024 to 28.6% by 2029. Studies initially highlighted the substantial benefits of ridesharing services for both the environment and the economy, including reduced pollution such as CO2 emissions and noise, decreased traffic congestion, and lower energy consumption [4]. Despite these promising advantages, concerns about climate change, traffic congestion, and oil dependency have emerged as ridesharing services have gained popularity [7]. For example, pickups and drop-offs from ridesharing services have increasingly disrupted traffic flow, particularly in urban areas that are already congested [11]. Additionally, other studies indicate that ridesharing services can raise energy consumption through deadhead miles driven without passengers which can negate any potential energy savings and exacerbate traffic congestion. In fact, a rideshared car ride generates 47% more CO2 emissions on average compared to an equivalent private car ride [18, 30].
During high-demand hours, ridesharing platforms face distinct challenges. The imbalance between available drivers and ride requests often leads to driver selectiveness, with drivers rejecting trips to less popular destinations or passengers with specific needs. This selectiveness can drive up trip cancellation rates, as platforms may prioritize less desirable trips in the interest of fairness. However, this puts pressure on drivers, who risk deactivation if they cancel too many rides, potentially exacerbating driver shortages as they leave the platform [22]. Even in centralized ridesharing systems where algorithms handle rider-driver assignment, online assignment systems often struggle to optimize objectives like minimizing wait times and reducing carbon emissions—particularly when driver availability is limited. In such cases, a simple assignment algorithm that assigns riders to the first available driver offers little flexibility in optimizing these factors. Consequently, much of the rider’s wait time is spent in the assignment queue, awaiting a match. Under these conditions, first-come, first-served strategies can result in extended deadhead miles and longer wait times, degrading the user experience for both riders and drivers.
@@
While some studies have explored methods to reduce deadhead miles and their associated emissions [18], and others have aimed to improve service quality by minimizing wait times [29, 30], many of these approaches lack performance guarantees across varying conditions. This issue becomes especially critical during high-demand periods when driver availability is low. In such cases, online algorithms without performance guarantees can fall significantly short compared to the optimal offline algorithms, which benefit from knowing future inputs. This underscores the need for new strategies that ensure environmentally sustainable ridesharing systems with reliable emissions reductions, even during peak demand, while also maintaining high service quality with respect to wait times.
@@
Developing such systems and strategies presents significant challenges. Providing theoretical guarantees in ridesharing optimization is inherently complex, as it requires managing dynamic decision-making processes involving both current and future drivers and riders. A major challenge lies in the uncertainty of future ride requests’ timing and locations. Assigning a driver to a rider not only affects the current ride but also determines when and where that driver will be available again, influencing the deadhead distance and emissions of subsequent rides. This creates a feedback loop, potentially causing long-term changes in the actions taken by the online algorithm and its overall performance. Furthermore, designing an online algorithm that balances the platform’s goals—such as emission reduction—with rider preferences, like minimizing wait times, adds another layer of complexity. Real-world factors, such as fluctuating urban traffic conditions, further complicate efforts to ensure consistent, optimal performance in practice.
@@
In this paper, we propose a Lyapunov-based Algorithm for Ride Assignment (LARA), designed to reduce carbon emissions in ridesharing platforms while minimizing rider wait times. While LARA is applicable at all times, its benefits over previous state-of-the-art algorithms are especially pronounced during high-demand hours, when a naive assignment approach could lead to long queues of unassigned requests and long waiting times accordingly. To address this, LARA dynamically adjusts the upper bound on deadhead distances for assigned drivers based on real-time conditions. The algorithm’s decisions are influenced by the number of ride requests in the assignment queue (i.e., unassigned requests waiting to be dispatched). We present a theoretical performance analysis of LARA and empirically evaluate it against state-of-the-art algorithms using both synthetic and real-world datasets.
Our contributions are summarized as follows:
  • We conduct a case study on the impact of limiting deadhead distances on both the carbon emissions of trips and rider waiting times. Based on this study, we propose a new approach to control the trade-off between emission and waiting times, framing it as a deadhead control problem (DCP). We model the objective of DCP as a weighted sum of two factors: the expected carbon reduction and the rate of ride assignment within the platform.
  • We introduce LARA, a near-optimal online algorithm for solving DCP that does not require future input predictions. LARA adjusts the upper bound on deadhead distances based on the number of ride requests in the assignment queue. It uses a tunable hyperparameter that allows for performance optimization under different conditions.
@@
  • We provide a theoretical performance analysis of LARA, showing that its objective value for DCP is within a bounded distance from the optimal solution (Theorem 1). We also show that this bound approaches zero when LARA is not constrained by the assignment queue length.
@@
  • We conduct a comprehensive experimental analysis to evaluate LARA using both real-world and synthetic datasets, comparing its performance against the existing emission-aware ride assignment algorithm TORA [30] as well as additional heuristic methods. Our results indicate significant reductions in both emissions and waiting times during high-demand periods. For instance, compared to classic approaches, LARA achieves up to $34.2\%$ reduction in average emissions on the synthetic dataset (Figure LABEL:fig:synthetic_batchDur) and up to $13.9\%$ on the real-world dataset (Figure 4), consistently outperforming competing algorithms across various scenarios. Furthermore, our analysis of LARA reveals a trade-off between reducing emissions and ensuring fair ride assignments among different drivers.
2 Case Study on Limiting Deadhead distances
The ride assignment problem extends the classic online asymmetric multiple traveling salesman problem, where $M$ traveling salesmen aim to visit $N$ cities such that the total travel distance is minimized [24, 19]. In our context, drivers take on the role of salesmen, and ride requests represent the cities. Formulating the ride assignment problem and developing an efficient algorithm is particularly challenging due to uncertainties about future ride requests, as well as the significant temporal and spatial dependencies between the decisions made by the online algorithms and the future state of the platform. Small changes in the current assignments can affect the set of available future drivers, and consequently the deadhead distances and emissions. Thus, rigorous formulation of this problem must account for the complex and time-coupled relationships between drivers and riders.
@@ 2 Case Study on Limiting Deadhead distances
The ridesharing platform consists of $M$ drivers and $N$ ride requests. Each ride request, indexed by $n$, includes a request time, a pickup location, and a drop-off location. Each driver, indexed by $m$, operates a vehicle that emits $e_{m}$ grams of CO2 per unit of distance and can serve only one ride at a time. Ride requests arrive sequentially, and upon each new request, the online assignment algorithm updates the locations and status of drivers before assigning a ride. Once assigned, the driver picks up the rider and is unavailable to serve other riders until the drop-off is complete. The objective is to minimize two key terms: 1) the total carbon emissions produced by the vehicles; and 2) the total waiting time for riders. Let $l^{\texttt{d}}_{n,m}$ represent the deadhead distance for driver $m$ to serve ride request $n$, and let $l^{\texttt{t}}_{n}$ represent the trip distance for request $n$. The emissions associated with serving ride request $n$ by driver $m$ can be expressed as:
$\text{Emission}_{n,m}=(l^{\texttt{d}}_{n,m}+l^{\texttt{t}}_{n})\cdot e_{m}.$ (1)
The emissions for each ride depend on both the deadhead and trip distances, as well as the vehicle’s emission rate. Another critical factor in the objective of this problem is the waiting time of the rider, which greatly influences user satisfaction and the overall quality of the ridesharing experience. Although shorter waiting times are highly desirable for riders, balancing minimal waiting times with optimizing emissions is complicated. Serving ride requests quickly while simultaneously minimizing emissions often requires trade-offs, making it difficult to achieve both objectives simultaneously.
Designing and analyzing an algorithm for online ride assignment problem presents significant challenges due to the intricate temporal and spatial dependencies between drivers and riders. The current state of the ridesharing platform, including the set of available drivers and their locations, directly impacts immediate outcomes like rider wait times and carbon emissions, while also shaping the platform’s future state. Even small adjustments in assignment decisions for individual rides can ripple through the system, affecting not just the current assignment but also future driver availability and locations. These seemingly minor changes can lead to substantial shifts in overall platform performance, influencing both environmental factors and service quality metrics. This interdependence creates a complex feedback loop, where present decisions have long-term repercussions, complicating the balance between short-term and long-term objectives in algorithm design. Due to these complexities, developing (near)-optimal algorithms for online ride assignment problem with rigorous theoretical understanding the algorithm’s performance is not tractable. As an alternative approach, and motivated by the following case study, in Section 3, we formulate another optimization problem with a more tractable objective function of controlling the deadhead distance.
2.1 Motivational Example
Addressing the original online ride assignment problem requires complex modeling and formulation, and finding a near-optimal solution involves intricate theoretical analysis. However, in what follows and through a preliminary case study, we show that alternative approaches can manage the trade-off between minimizing emissions and reducing waiting times in ridesharing platforms. More specifically, we present a case study that examines the impact of limiting deadhead distances on average emissions and rider waiting times. The insights gained from this study inspire a new approach to balancing these two critical performance metrics in ridesharing services. This study focuses on minimizing total carbon emissions while maintaining reasonable waiting times for riders. We evaluate the performance of an online ride-assignment algorithm designed to balance these objectives.
To implement our case study, we utilize real-world ride request data from the RideAustin dataset [25] in combination with a parametric online ride-assignment algorithm. This dataset provides comprehensive data for simulating diverse ridesharing scenarios, allowing us to evaluate both average waiting times and emissions under various operating conditions. More details about the RideAustin dataset are given in Section 5.2.
The ride-assignment process in our study operates in a batch system, where the algorithm attempts to assign unallocated ride requests to available drivers every two minutes. If a request remains unassigned during a batch cycle, it is retained in the queue for the next iteration. This batch-based approach mirrors common industry practices, where the ride assignments are performed at regular intervals to strike a balance between computational efficiency and service responsiveness.
We present a simple yet effective algorithm, where the goal is to minimize carbon emissions by limiting the deadhead distance within a predefined threshold. This threshold plays a critical role in managing deadhead miles, directly impacting both rider waiting times and total emissions. The parametric algorithm uses a fixed distance threshold, denoted as $d$, within which it attempts to assign drivers to passengers. The objective is to minimize emissions for each trip while ensuring that the deadhead distance does not exceed $d$. If no driver is available within this threshold, the ride request is deferred to the next batch cycle for reassignment. For our evaluation, we focused on ride requests from December 1, 2016, in the RideAustin dataset, which includes 2,909 ride request traces.
Figure LABEL:fig:motivation_plots presents the results of our experiments, showing the average waiting time and average carbon emissions per trip for different values of the threshold distance $d$. As expected, increasing the threshold $d$ enhances the likelihood of finding a driver with lower per-mile carbon emissions. This allows the online assignment algorithm to reduce the total carbon emissions for each trip by pairing passengers with more efficient drivers, even if they are further away. However, this comes at the cost of increased expected deadhead distances, which in turn leads to longer waiting times for passengers. Increasing the deadhead distance limit from $5$ to $50$ km results in $33\%$ reduction in the average emission of trips while the average waiting time of riders increases by $52\%$. The results clearly illustrate the significant trade-off between minimizing average waiting time and reducing carbon emissions—a key consideration for optimizing ridesharing services.
These findings demonstrate that an intelligent online algorithm has the potential to effectively manage both average waiting time and carbon emissions by dynamically adjusting the threshold value $d$ based on current conditions. During peak demand periods, when there is a long queue of unassigned ride requests, a substantial portion of the passenger’s waiting time occurs while they wait in the queue for their turn to be assigned. In such situations, the algorithm should prioritize reducing waiting times by selecting lower values for $d$, thereby increasing the frequency of ride assignments. This approach helps maintain a high level of service responsiveness, which is critical during times of high demand. While this approach may seem to favor drivers with low-emission vehicles by assigning them more ride requests, our experimental analysis disproves this hypothesis (See Section 5.2).
Conversely, during off-peak periods, when the number of unassigned ride requests is low, the algorithm can afford to prioritize environmental considerations by opting for higher values of $d$. This allows the system to minimize the emissions of each trip by assigning passengers to the more efficient drivers, even if they are located further away. This dynamic adjustment of $d$ is a challenging task for an online algorithm, particularly because of the inherent uncertainty in predicting future ride requests. Factors such as the rate of new ride requests and the average trip distances may fluctuate over time, complicating the algorithm’s ability to adapt to changing patterns.
For instance, during special events or unexpected surges in demand, the patterns of ride requests may deviate significantly from the norm, potentially leading to suboptimal performance by a naive online algorithm. This highlights the need for more rigorous algorithms capable of learning and adapting to the evolving environment, potentially incorporating predictive modeling or machine learning techniques to better anticipate future ride request patterns.
The results and discussion presented here propose an alternative approach for managing emissions and wait times during high-demand hours, motivating the development of an online algorithm that dynamically adjusts the deadhead distance limit to minimize both average emissions and rider wait times in such hours. In the next section, we formally model the problem of controlling deadhead distances in an online setting. Then, in Section 4, we present a near-optimal algorithm that tackles the deadhead control problem in an online manner and analyze its theoretical performance.
3 Online Deadhead Control Problem
To address the challenges associated with the design of online algorithms for emission-aware ride assignment problem, and motivated by the example in Section 2.1, in this section, we formulate the online Deadhead Control Problem, referred to as DCP, as an alternative approach to controlling emissions and deadhead distances within ridesharing platforms. In DCP, the online algorithm’s task is to determine the maximum acceptable deadhead distance for rides assigned in each upcoming batch of assignments. The objective of DCP is adjusted for the traceability of theoretical analysis but still focuses on minimizing both total emissions and passenger wait times within the ridesharing system.
In the DCP, the goal is to select an upper bound for the deadhead distance of rides in the next assignment batch. Each ride request is then assigned to a driver within this bound, aiming to minimize the overall emissions for the assigned rides. This approach represents an adjustable version of the algorithm used in Section 2.1. Algorithm 1 outlines the pseudocode for the ride assignment process in DCP. The objective function in DCP consists of two components: $E_{N}$, which represents the time-averaged expected reduction, and $S_{N}$, which represents the rate of ride assignments (the average number of assignments completed over a fixed period of time on the platform). The first term, $E_{N}$, captures the emissions aspect, and the term $S_{N}$ focuses on reducing passenger wait times.

### v2

@@
Ridesharing has experienced significant global growth over the past decade and is becoming an integral component of modern transportation systems. However, despite their benefits, ridesharing platforms face fundamental inefficiencies that contribute to negative environmental impacts. A prominent source of such inefficiency is the “deadhead miles”—the distance traveled by vehicles without passengers between trips—which accounts for a substantial portion of carbon emissions. This issue becomes especially severe during high-demand periods, when the volume of ride requests exceeds the available driver supply, leading to suboptimal rider-to-driver assignments, longer deadhead trips, and increased emissions. Although limiting these unproductive miles can reduce emissions, doing so may increase passenger wait times due to limited driver availability, thereby degrading the overall service experience. As ridesharing continues to scale, there is a critical need for environment-aware solutions that jointly minimize emissions and maintain high-quality service, particularly in terms of rider wait times.
In this paper, we introduce LARA, an online rider-to-driver assignment algorithm that dynamically adjusts the maximum allowable distance between rider and drivers and assigns ride requests accordingly. While LARA is applicable in general settings, it is particularly effective during peak demand periods, achieving reductions in both emissions and wait times. We provide theoretical guarantees showing that LARA achieves near-optimal performance in online environments, with respect to an optimal offline benchmark. Beside our theoretical analysis, our empirical evaluations on both synthetic and real-world datasets show that LARA achieves up to a $34\%$ reduction in carbon emissions and up to a $50\%$ decrease in rider wait times, compared to state-of-the-art baselines. While prior work has explored emission-aware ride assignment, LARA is, to our knowledge, the first algorithm to offer both rigorous theoretical guarantees and strong empirical performance.
@@ 1 Introduction
Ridesharing services have revolutionized urban mobility by providing convenient, on-demand transportation and have seen widespread global adoption [40]. The global ridesharing market is projected to reach \$212B by 2029, with an estimated 2.3B users [35], representing a 167% increase in market size and a 53% increase in user base compared to 2020. Although these services were initially perceived as environmentally friendly—promising reductions in emissions and congestion—recent studies have shown otherwise. A key contributor to the increased emissions is deadheading: miles driven by drivers without passengers. These deadhead miles can significantly increase overall energy consumption, and in fact, a rideshared trip has been shown to generate approximately 47% more CO2 emissions than a comparable private car trip [18, 33, 27, 34]. Given their scale and growing influence on modern transportation systems, understanding and mitigating the environmental impacts of ridesharing has become increasingly critical.
During peak demand hours, ridesharing platforms face challenges like driver selectiveness and trip cancellations, leading to assignment inefficiencies [24]. Even centralized assignment algorithms struggle to optimize both wait times and emissions when driver availability is low. Simple assignment strategies, such as first-come, first-served, can increase deadhead miles and worsen user experience. Addressing these inefficiencies requires intelligent online algorithms that balance service quality and sustainability. Moreover, accurately predicting the exact timing and magnitude of peak demand introduces significant uncertainty, which can further reduce the effectiveness of even advanced assignment strategies [12, 17, 16].
@@
While recent studies have explored methods to reduce deadhead miles and their associated emissions [18], and others have aimed to improve service quality by minimizing wait times [32, 33, 34], many of these approaches lack performance guarantees across varying conditions. This issue becomes especially critical during high-demand periods when driver availability is low. In such cases, online algorithms without performance guarantees can fall significantly short compared to the optimal offline algorithms, which benefit from knowing future inputs. This underscores the need for new strategies that ensure environmentally sustainable ridesharing systems with reliable emissions reductions, even during peak demand, while also maintaining high service quality with respect to wait times.
Figure 1: Adjusting the maximum allowable deadhead distance could balance the trade-off between average emission and rider waiting times within the ridesharing system.
@@
Developing such strategies with theoretical guarantees in ridesharing optimization is inherently complex, as it requires managing dynamic decision-making processes involving both current and future drivers and riders. More specifically, a major challenge lies in the uncertainty of future ride requests’ timing and locations. Assigning a driver to a rider not only affects the current ride but also determines when and where that driver will be available again, influencing the deadhead distance and emissions of subsequent rides. This causes long-term changes in the online algorithm’s actions and overall performance. Furthermore, designing an online algorithm that balances the platform’s sustainability goals—such as emission reduction—with rider preferences, like minimizing wait times, adds another layer of complexity. Real-world factors, such as fluctuating urban traffic conditions, further complicate efforts to ensure consistent, optimal performance in practice.
Contributions. Due to the above complexities, formulating a tractable multi-objective ride assignment problem that lends itself to rigorous algorithmic design and theoretical analysis poses significant challenges. As an alternative, inspired by the empirical findings presented in Section 2 (see Figure 1), we propose an alternative approach to control the trade-off between emissions and waiting times, framing it as a Deadhead Control Problem (DCP). We model the objective of DCP as a weighted sum of two factors: the expected carbon reduction and the rate of ride assignment which directly impacts the rider waiting time. As a ride assignment problem, this approach implicitly balances the trade-off between emissions reduction and rider wait times and, more importantly, is tractable for rigorous algorithm design and performance analysis.
@@
In Section 3, we solve DCP by proposing a Lyapunov-based Algorithm for Ride Assignment (LARA), an online algorithm designed to reduce carbon emissions in ridesharing platforms while minimizing rider wait times. LARA’s benefits over previous state-of-the-art algorithms are especially pronounced during high-demand hours, when a naive assignment approach could lead to long queues of unassigned requests leading to long waiting times. LARA dynamically adjusts the upper bound on deadhead distances for assigned drivers based on real-time conditions. The algorithm’s decisions are influenced by the number of ride requests in the assignment queue (i.e., unassigned requests waiting to be dispatched), and it uses a tunable hyperparameter that allows for performance optimization under such dynamic conditions.
@@
In Section 4, we analyze the theoretical performance of LARA. We show that the objective value obtained by LARA for DCP is within a bounded distance from the optimal solution when extra (Theorem 1) and no additional (Theorem 2) information about environment is provided in advance. We also show that this bound approaches zero when LARA is not constrained by the assignment queue length, which essentially shows the optimality of the algorithms under high-demand and rush hour periods.
@@
Finally, in Section 5, we empirically evaluate the performance of LARA using both synthetic and real-world datasets. Comparing its performance against the existing emission-aware ride assignment algorithm TORA [33] as well as additional heuristic methods. Our results indicate significant reductions in both emissions and waiting times during high-demand periods. For instance, compared to state-of-the-art emission and waiting time aware baselines, LARA achieves up to $34\%$ reduction in average emissions on the synthetic dataset (Figure LABEL:fig:synthetic_batchDur) and up to $13.9\%$ on the real-world dataset (Figure 4), consistently outperforming competing algorithms across various scenarios. Furthermore, our analysis of LARA reveals a trade-off between reducing emissions and ensuring fair ride assignments among different drivers. This trade-off in the algorithm design could be leveraged by the ridesharing platform to balance between sustainability and customer satisfaction goals in runtime.
2 Problem Formulation
@@
The ridesharing platform consists of $M$ drivers and $N$ ride requests, where $N$ is unknown in advance and ride requests arrive sequentially over time. Each ride request, indexed by $n$, includes a request time, a pickup location, and a drop-off location. Each driver, indexed by $m$, operates a vehicle that emits $e_{m}$ grams of CO2 per unit distance and can serve only one ride at a time. Ride requests arrive sequentially, and upon each new request, the online assignment algorithm updates the drivers’ locations and statuses before assigning a ride. Once assigned, a driver picks up the rider and remains unavailable until the drop-off is complete.
The objective is to minimize two key factors: (1) total carbon emissions and (2) average rider waiting time. Carbon emissions for a trip depend on both the deadhead distance and the trip distance, as well as the vehicle’s emission efficiency. Rider waiting time consists of the time period the rider waits until the assignment happens and the time until the driver arrives at the pickup location. While shorter waiting times benefit riders, minimizing both waiting time and emissions requires trade-offs, making it challenging to optimize both simultaneously.
Previous studies [33, 32, 34] highlight the trade-off between minimizing deadhead emissions and reducing passenger wait times in ridesharing platforms. Assigning passengers to drivers with low-emission vehicles can reduce emissions but may increase wait times due to longer pickup distances. Conversely, imposing a strict limit on the maximum allowable distance between passengers and assigned drivers can shorten wait times but reduces the chances of finding nearby low-emission vehicles, potentially increasing overall emissions. To better capture the trade-off between emission and waiting times in the ridesharing platform, we performed a simple analysis over the ride requests posted during first week of December 2016 from RideAustin dataset [28] (see Section 5 for details). We imposed a cap on the deadhead distance of eligible drivers and assigned each rider to the driver with the lowest emissions. Then, we analyzed the impact of this limit on deadhead distances on the performance of ridesharing system. The results in Figure 1 indicate that instead of directly modifying ride assignment strategies based on platform status, the online algorithm can adjust the maximum allowable deadhead distance for assigned rides, balancing the trade-off between average emissions and rider waiting times. This insight motivates an alternative approach to modeling ridesharing optimization with a more tractable theoretical analysis. In the following, we introduce a novel problem formulation that captures the objectives and optimization challenges of ridesharing systems.
2.1 DCP: Deadhead Control Problem
Motivated by the insights from our preliminary empirical results in Figure 1, and to address the challenges of online algorithm design for emission-aware ride assignment, we introduce the Online Deadhead Control Problem, referred to as DCP. This problem offers a new perspective on managing emissions and deadhead distances in ridesharing platforms. In DCP, ride assignments are made in batches, and the algorithm determines the maximum allowable deadhead distance for each upcoming batch. The objective is to jointly minimize total emissions and passenger wait times over the entire time horizon. Each ride request is then matched to a driver within this limit, with a focus on reducing emissions. The resulting online ride assignment process is detailed in Algorithm 1. In line 2, an online deadhead control algorithm determines the deadhead limit to be used during the next batch assignment. This limit, in turn, defines the set of eligible drivers for each ride request in line 4.